OpenStudy (anonymous):

The Big Screamer Coaster carries 92 people altogether. Some of its cars carry 4 passengers and the rest carry 6 passengers. There are 3 less 6-passenger cars the there are 4-passenger cars. Write and equation to solve how many 4-passenger cars there are.

OpenStudy (anonymous):

set up a system of equation. let the six passenger cars be s and the four passenger cars be f. the number of total people is 92 so you can make your first equation. 6s + 4f = 92. Now onto the second equation, since there are 3 less six passenger cars than four passenger cars, you can make the second equation to be f - s =6 so your system is 6s + 4f = 92 f - s = 6 now solve the system for f